RC meeting and Run Plan, Wednesday, January 18th
From SBS wiki
Return to the RC Meeting Calendar
RC Meeting Information
|Current RC:||Donald Jones|
|Time:||Daily at 3:30PM|
|Location:||Counting House Conference Room|
Remote Connection Details
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
- Beam is expected back into the hall late swing. SL follow beam checkout procedures . TO make sure that after the SBS and BB magnets are turned ON, the fields are rotated into the SBS/BB ON/ON configuration and perform an NMR measurement.
- RF work will continue tomorrow to optimize high and stable beam current to both Halls A & C.
Current Status and Run Plan
- Provakar looked into the noisy channels on BBCAL Shower 7.7 and 21.7 on the scope. They looked just fine with SBS/BB OFF HV configuration. They seem to get noisier due to the fringe fields .
- Jiwan looking into hot channels (channel 74) and FIFO readout issue (update?)
- Addressing low efficiency strip on BB GEM layer 2 - easy fix unsuccessful (update?)
- LVPS replaced for SBS side GEMs 
- Timing Hodoscope:
- Ralph fixed the hodoscope channel 2R issue and moved it to another HV supply which isn't currently read out by EPICS. The current slot on the HV GUI (which used to be 2R) is now empty and stays at -500V. Will look into integrating this channel into EPICS data stream later.
- Moller power supply work
- Work done in BSY yesterday and today morning. PS stable at 430A and is enough to schedule a Moller measurement. Revisit the question on when to do a Moller measurement on Monday.
- 'Heavy gas' bottle replacement completed by Carlos and Todd with Jack's support.
- Parity DAQ:
- Charge asymmetry for parity runs 10423 through 10441 (Jan 17 12:57 - Jan 18 08:07) .
- Reference cell:
- Todd restarted the turbo pump 
- Oven temperature increased by 5 deg. Now at 260C. Now that we have established high current on the polhe3 cell, need to focus efforts on optimizing the polarization.
- Reference cell:
- Optimize target polarization
Please consider signing up for shifts .